# Anyon Quasilocalization in a Quasicrystalline Toric Code

### v1

@@ 55 $(35J^{5}/128J_{z}^{4})$ 0
Table 3: The list of $A_{s}^{(n)}$ and $B^{(n)}_{p}$ coupling parameters of the toric code used in the main text. We consider the case: $J_{x}=J_{y}=J\ (\ll J_{z})$. See Fig.S4 for the distribution of these couplings for a specific generation.
@@ IV Energy spectra for spatially non-uniform and uniform $\lambda_{A}^{(n)}$
See Fig.S6 for the low-energy single–ee–anyon spectra in two representative cases: (1) the Gen-7 lattice with spatially non-uniform onsite potentials $\lambda_{A}^{(n)}$ as listed in Table.3, and (2) the Gen-5 lattice with uniform onsite potentials $\lambda_{A}^{(n)}=\lambda_{A}^{(2)}$ for all $n$, and in presence of the background $\pi$-flux distribution.
Figure S6: The $e$ anyon energy spectra: (a) Energy eigenvalues of the effective low-energy bosonic hopping model [Eq.(13) of main text] with non-uniform onsite potentials $\lambda_{A}^{(n)}$ are plotted as a function of the Zeeman coupling $h$ for the Gen-7 lattice (1906 sites). The characteristic four-band structure—exponentially separated in energy—and both linear ($\sim h$) and quadratic ($\sim h^{2}$) scaling of the levels, observed earlier for the Gen-5 lattice (see main text), persist here. (b) The spectrum in the absence of onsite potential non-uniformity ($\lambda_{A}^{(n)}=\lambda_{A}^{(2)}$ for all $n$) is shown as a function of $h$. In this case, all energy levels exhibit linear scaling with $h$. The Gen-7 spectrum shows a similar behavior, but with a denser distribution of levels owing to the larger system size.
@@ VI Dependence of plateau residence time on $J/J_{z}$
We initialize the wavepacket at an $A^{(3)}$-type vertex, which belongs to the class of vertices connected by $O(h)$ hopping amplitudes. Its time-evolved probability distribution is shown in Fig.4(d) of the main text. Our goal is to determine the time scale required for the wavepacket to spread from the smaller-radius ring to the larger-radius ring. As shown in Fig.S10(a), the plateau width (or equivalently $T_{p}$) varies only weakly with $g$ over a broad range. However, the behavior changes drastically near $g=1$, where, as shown in Fig.S10(b), $T_{p}$ drops sharply. The full functional dependence on $g$ is shown in panel (c).

### v2

@@ 55 $(35J^{5}/128J_{z}^{4})$ 0
Table 3: The list of $A_{v}^{(n)}$ and $B^{(n)}_{p}$ coupling parameters of the toric code used in the main text. We consider the case: $J_{x}=J_{y}=J\ (\ll J_{z})$. See Fig.S4 for the distribution of these couplings for a specific generation.
@@ IV Energy spectra for spatially non-uniform and uniform $\lambda_{A}^{(n)}$
See Fig.S6 for the low-energy single–ee–anyon spectra in two representative cases: (1) the Gen-7 lattice with spatially non-uniform onsite potentials $\lambda_{A}^{(n)}$ (see Eq.(14) of main text), and (2) the Gen-5 lattice with uniform onsite potentials $\lambda_{A}^{(n)}=\lambda_{A}^{(2)}$ for all $n$, and in presence of the background $\pi$-flux distribution.
Figure S6: The $e$ anyon energy spectra: (a) Energy eigenvalues of the effective low-energy bosonic hopping model [Eq.(14) of main text] with non-uniform onsite potentials $\lambda_{A}^{(n)}$ are plotted as a function of the Zeeman coupling $h$ for the Gen-7 lattice (1906 sites). The characteristic four-band structure—exponentially separated in energy—and both linear ($\sim h$) and quadratic ($\sim h^{2}$) scaling of the levels, observed earlier for the Gen-5 lattice (see main text), persist here. (b) The spectrum in the absence of onsite potential non-uniformity ($\lambda_{A}^{(n)}=\lambda_{A}^{(2)}$ for all $n$) is shown as a function of $h$. In this case, all energy levels exhibit linear scaling ($\sim h$). The Gen-7 spectrum shows a similar behavior, but with a denser distribution of levels owing to the larger system size.
@@ VI Dependence of plateau residence time on $J/J_{z}$
We initialize the wavepacket at an $A^{(3)}$-type vertex, which belongs to the class of vertices connected by $O(h)$ hopping amplitudes. Its time-evolved probability distribution is shown in Fig.4(d) of the main text. Our goal is to determine the time scale required for the wavepacket to spread from the smaller-radius ring to the larger-radius ring. As shown in Fig.S10(a), the first plateau width (or equivalently $T_{p}$) varies only weakly with $g$ over a broad range. However, the behavior changes drastically near $g=1$, where, as shown in Fig.S10(b), $T_{p}$ drops sharply. The full functional dependence on $g$ is shown in panel (c).